The sum of 3 consecutive integers is 114. find the numbers
Answer by ewatrrr(10682)


Hi,
Let x , (x+1) and (x+2) represent the 3 consecutive integers
Question States***
x + (x+1) + (x+2)= 114
Solving for x
3x + 3 = 114
3x = 111
x = 37 Integers are 37, 38, 39
CHECKING our Answer
37 + 38 + 39 = 114